OpenStudy (anonymous):

In triangle ABC, if a = 10, b = 7, and c = 8, the the value of cosC is..

OpenStudy (eric_d):

Use the law of cosines c^2 = a^2 + b^2 - 2ab cos C

OpenStudy (eric_d):

Solve this: 64=100+49-140cos C

OpenStudy (wolf1728):

See attached C = 52.617° So cosine C = 0.60714